a student pushed a box 16.0 m across a smooth, horizontal floor using a constant force of 124 N. If the for applied for 8.00 s, how much power did it take to accomplish this task ( to the nearest watt)?
1
Expert's answer
2020-02-03T14:06:58-0500
"W = \frac{F * S}{t}= \frac{124*16}{8}= 248 W"